OpenStudy (babyslapmafro):

I found the derivative of a functino and now I need help finding the critical/stationary values. (click to see)

OpenStudy (babyslapmafro):

\[ f'(x)=\frac{ x^2+2x-1 }{ x^2+2x+1 }\]

OpenStudy (anonymous):

Bottom can be factored to find vertical asymptotes. Top looks to require quadratic formula to find zeroes.

OpenStudy (babyslapmafro):

ok thanks

OpenStudy (anonymous):

If you want stationary inflection points, you'll need to find where the second derivative =0.

OpenStudy (babyslapmafro):

I just needed to find the abs max/min

OpenStudy (babyslapmafro):

I found that there is only an absolute max within the interval I am trying to find. The x-value for the abs max is -2-sqrt2

OpenStudy (babyslapmafro):

How do I find an exact value for the y-coordinate of the abs max? Where the original function is...

OpenStudy (babyslapmafro):

\[ f(x)=\frac{ x^2+1}{ x+1}\]

OpenStudy (anonymous):

If you have the original function, plug-and-chug.

OpenStudy (babyslapmafro):

I have found a decimal approximation but I do not know how to obtain the exact value for y.

OpenStudy (anonymous):

Hmm. Gimme a sec to work it out, I'll let you know what I see.

OpenStudy (anonymous):

What is the interval you're looking in?

OpenStudy (babyslapmafro):

(-5,-1)

OpenStudy (anonymous):

I'm getting \(-1-\sqrt{2}\) for the critical point.

OpenStudy (babyslapmafro):

ya i just checked the book, and they have the same value...

OpenStudy (babyslapmafro):

i just realized i got -2-sqrt because i only divided 2 from the coefficient of sqrt2 and not the integer -2

OpenStudy (anonymous):

If you want the exact y value, then you need to put the exact x value in the function. \[f(-1-\sqrt{2})=\frac{(-1-\sqrt{2})^2+1}{-\sqrt{2}}\]

OpenStudy (anonymous):

Also, you'll either need to look for intervals of increase and decrease, or put the critical point into the second derivative to test for concavity to determine rel. min. or max.

OpenStudy (babyslapmafro):

Yes, I used a number line and found the values for f'(x)

OpenStudy (anonymous):

Ok, so to verify absolute extrema, you also need to check the endpoints of the interval.

OpenStudy (babyslapmafro):

Even if it is an open interval?

OpenStudy (babyslapmafro):

@CliffSedge

OpenStudy (anonymous):

I'm pretty sure, yeah. Even if it's open, you can find the limit of the function at those points.

OpenStudy (babyslapmafro):

From Wolframalpha: "An open interval is an interval that does not include its end-points."

OpenStudy (anonymous):

There is that asymptote at -1, though.

OpenStudy (anonymous):

Right, the function is not defined at the endpoints, but the function still exists in that neighborhood.

OpenStudy (anonymous):

in other words. You wouldn't bother checking -5, but you could check -4.99999999999

OpenStudy (babyslapmafro):

That said, to find the min/max near the end-points I believe I would have to choose a value sufficiently close to -5 and -1 but where and how would I determine what is close enough?

OpenStudy (anonymous):

Use properties of limits. You should be able to use -5 as is no problem. -1 is a discontinuity though, so you'll need something like L'Hopital's rule to find the limit there.

OpenStudy (babyslapmafro):

Ok, well thank you for the help. I will find out if I need to find the limits close to the ends of an open interval from my teacher just to be sure. Thanks a lot for the help...

OpenStudy (babyslapmafro):

excuse me, not the limits, the max/min values...

OpenStudy (anonymous):

When in doubt, graph it.